OpenStudy (rayman5766):

Three friends entered a contest to count the number of jelly beans in a jar. The table shows their guesses of the number of jelly beans in the jar. There were actually 950 jelly beans in the jar. Name Number guessed Brandon 991 Li Mei 939 Jodi 976 A. Li Mei’s estimate had about a 1.2% error. B. Jodi had the highest percent error. C. Jodi’s estimate had about a 26% error. D. Brandon’s estimate had about a 4.3% error

OpenStudy (rayman5766):

@sleepyhead314 help me also I have a question for you

OpenStudy (anonymous):

What is your question????

OpenStudy (sleepyhead314):

so try calculating the percent error for each person :) can you do that?

OpenStudy (rayman5766):

dud I am sorry but its so lagy

OpenStudy (rayman5766):

a and c

OpenStudy (anonymous):

@rayman5766 no bro its fine just wondering cuz i was wondering

OpenStudy (rayman5766):

oh thanks for understanding

OpenStudy (sleepyhead314):

neither a nor c

OpenStudy (anonymous):

so its either b or d

OpenStudy (rayman5766):

I think I under stand its b

OpenStudy (sleepyhead314):

*cries*

OpenStudy (sleepyhead314):

brandon's percent error = (991-950) / 950 * 100 = ?

OpenStudy (rayman5766):

its b and d thanks it says 2 answers

OpenStudy (sleepyhead314):

oh then not b and Li Mei's percent error = (950-939)/950 * 100 = ?

OpenStudy (sleepyhead314):

brandon's percent error = (991-950) / 950 * 100 = ? and Li Mei's percent error = (950-939)/950 * 100 = ? should get you close to the answers you need

OpenStudy (rayman5766):

oh i did my cousin helped me thanks he said it is b and c